OpenStudy (anonymous):

6x/11 + 9/2 = 3 what is the LCD

OpenStudy (anonymous):

so start off by finding x , 3- 9/2 = -1.5 so 6x/11 = -1.5 hence, x = -2.75 6*(-2.75) / 11 = -1/4 so i guess 4 would be the lowest common denominator since 9/2 can be changed into 18/4

OpenStudy (kaylee_td):

LCD- the lowest common multiple of the denominators of several fractions

OpenStudy (kaylee_td):

(And welcome to OS \(\color{green}{\huge ツ}\) )